Simplify each of the following by hand. Show your working and give the
answers in the form a + bi (a, b are real).
(a) i^42
(b) 3+2i/2-5i
(c) (2 cos pi/12 + 2i sin pi/12)^4

Answer by tommyt3rd(5050)

I will do a)
i^42=(i^40)(i^2)=(1)(-1)=-1